OpenStudy (anonymous):

find two geometric means between 486 and 6

OpenStudy (primeralph):

Multiply them and take the square root of the result.

hartnn (hartnn):

if b is the geometric mean between a and c , then \(\large b^2 = ac \\ \large b = \sqrt {ac}\)

OpenStudy (anonymous):

let G1 and G2 be two G>means between 486 and 6 then 6 , G1 , G2 , 486 are in G.P 1st term = 6 last or 6th term = 486 and n =4 G1 =a*r G2 = a*r^2 you have already first term you have to just find common ratio 'r'